# OmniNWM: Omniscient Driving Navigation World Models

### v2

@@ Abstract
Autonomous driving world models are expected to work effectively across three core dimensions: state, action, and reward. Existing models, however, are typically restricted to limited state modalities, short video sequences, imprecise action control, and a lack of reward awareness. In this paper, we introduce OmniNWM, an omniscient panoramic navigation world model that addresses all three dimensions within a unified framework. For state, OmniNWM jointly generates panoramic videos of RGB, semantics, metric depth, and 3D occupancy. A flexible forcing strategy enables high-quality long-horizon auto-regressive generation. For action, we introduce a normalized panoramic Plücker ray-map representation that encodes input trajectories into pixel-level signals, enabling highly precise and generalizable control over panoramic video generation. Regarding reward, we move beyond learning reward functions with external image-based models: instead, we leverage the generated 3D occupancy to directly define rule-based dense rewards for driving compliance and safety. Extensive experiments demonstrate that OmniNWM achieves state-of-the-art performance in video generation, control accuracy, and long-horizon stability, while providing a reliable closed-loop evaluation framework through occupancy-grounded rewards.
@@ 1 Introduction
OmniNWM (Ours) R,S,D,O 321 Normalized Plücker Ray-map-based Integrated Occupancy-based
@@
  • Leveraging jointly generated 3D occupancy, OmniNWM naturally supports closed-loop navigation and evaluation through integrated, occupancy-grounded dense rewards.
@@
Figure 2: Architecture overview of OmniNWM. Our model takes as input a historical trajectory, a reference image, and flexible multi-level noise applied view-wise and frame-wise. It simultaneously forecasts panoramic RGB, semantic, and metric depth videos (Section 3.1.1), along with 3D occupancy (Section 3.1.2) and a future planning trajectory (Section 3.1.3). Note that input trajectories guide the generation of the current video clip, while future trajectories are forecasted for the next clip, forming a closed-loop system. The input trajectories with camera poses are encoded into normalized panoramic Plücker ray-maps, providing a unified pixel-level description and enabling precise action control (Section 3.2). A flexible forcing strategy injects independent multi-level noise at the view and frame levels to support robust and flexible auto-regressive generation (Section 3.3)
@@ 2 Related Work
World Models for Autonomous Driving. Recent progress in driving world models has advanced visual realism and controllability [94, 18, 40, 16, 42, 73, 8, 92, 84, 19, 93, 60, 103, 53, 43, 70]. Early efforts like DriveDreamer [72] leverage multi-stage pipelines to incorporate traffic rules. MagicDrive [17] introduces cross-view attention for street-view synthesis, while Vista [18] improves generation via latent replacement strategies. UniScene [43] proposes a hierarchical occupancy-centric framework. Other notable works include ReSim [87] for simulation data augmentation, Drive-WM [73] for multi-view forecasting, and GEM [23] for egocentric prediction. However, existing models typically lack unified multi-modal generation (e.g., semantic map, 3D occupancy)  [72, 17, 18] and integrated reward modeling for closed-loop evaluation [102, 16, 43]. Our work addresses these limitations with long-horizon, multi-modal generation, precise control, and integrated reward modeling.
Camera-controlled Video Generation. Camera control has become a critical technique for guiding generative models in dynamic 3D environments [24, 25, 75, 37, 4, 89, 29]. CameraCtrl [24] introduces a plug-and-play module to parameterize camera trajectories. MotionCtrl [75] decouples camera and object motion for independent control, while CamCo [83] introduces epipolar constraints to improve geometric consistency. VD3D [1] embeds Plücker coordinates into transformer architectures for view synthesis, and ReCamMaster [4] focuses on re-rendering existing videos from novel camera paths. However, these methods focus on short-range, monocular sequences, limiting their applicability to large-scale multi-view environments in autonomous driving [24, 25, 75, 83, 1, 4]. OmniNWM extends camera controllability to panoramic, long-horizon settings through a normalized Plücker ray-map representation.
@@ 3 Methodology
At its core, OmniNWM unifies the state–action–reward triad within a single framework, enabling comprehensive state generation, precise action control, and integrated dense rewards. As illustrated in Figure 2, OmniNWM supports multi-modal generation (Section 3.1), producing pixel-aligned panoramic videos (Section 3.1.1), as well as 3D semantic occupancy (Section 3.1.2) and future planning trajectories (Section 3.1.3). Section 3.2 introduces the normalized panoramic Plücker ray-map camera control scheme, while Section 3.3 describes the flexible forcing strategy for long-term robust generation.
@@ 3.1 Comprehensive Generation within OmniNWM
The overview of the model architecture is illustrated in Figure 2. We leverage a pretrained 3D VAE [36] to encode input video frames into compact spatiotemporal latents with a compression ratio of $4\times 8\times 8$. A panoramic diffusion transformer (PDiT) with cross-view attention layers [79, 44] is leveraged to denoise the resulting latents. The denoised latents are decoded into pixel-aligned RGB, semantic, and depth outputs, which are subsequently used to produce 3D semantic occupancy and future planning trajectories.
@@ 3.1.1 Pixel-level Aligned Multi-model Generation
where $\otimes$ denotes the outer product. $\texttt{Adap}_{d}$ and $\texttt{Adap}_{s}$ are adaptive aggregation networks for depth and semantic contexts. The resulting volume feature $\widehat{\mathbf{V}}$ is passed to a generation head with up-sampling and a softmax layer to produce the final 3D semantic occupancy. Unlike volumetric-based models like [43, 105] that directly synthesize full occupancy grids at high computational cost, our design computes occupancy from lightweight pixel-aligned features, which scales to long sequences and, crucially, enables integrated occupancy-grounded dense driving rewards.
@@ 3.1.3 Trajectory Planning for Closed-loop Navigation
As illustrated in Figure 1(c), OmniNWM functions as an environment simulator to enable closed-loop navigation with planning agents. Planning trajectories—generated from the model’s own outputs—guide future state generation when encoded into Plücker ray-maps.
@@ 3.2.2 Scale and Pose Invariant Normalization
As shown in Figure 5 (a), we apply scale and pose normalization to each frame ${P}_{j}\in\mathbb{R}^{6\times 6\times h\times w}$ of the panoramic ray-map to obtain a unified Plücker representation. Our normalization method decouples camera parameters: a shared intrinsic matrix from a reference camera (defaulting to the front view) ensures scale consistency, while individual extrinsics retain distinct viewpoints. These parameters are then transformed into a unified coordinate system. For each pixel $(u,v)$ in the $k^{\text{th}}$ camera, the direction vector is computed using the reference intrinsics ${K}_{0}$ and the camera’s rotation matrix ${R}_{k}$:
@@ 3.3 Long-term Generation with Flexible Forcing
As shown in Figure 6 (b), this formulation enables two distinct auto-regressive inference modes: (1) Frame-level auto-regression generates a single future frame conditioned on multiple preceding frames:
@@
where $f_{\theta}$ is the diffusion model. $K$ is the input context window. $\mathbf{\tilde{x}}^{(t)}$ is the potentially corrupted latent at time $t$. This mode is particularly suitable for frame-wise trajectory planning simulation (Figure 4). (2) Clip-level auto-regression generates multiple future frames from fewer inputs:
@@ 4.1 Experimental Setup
We train our model on the NuScenes [6] and NuScenes-Occupancy [71] datasets. Our Panoramic Diffusion Transformer (PDiT) comprises 11.22B parameters, including an 11B Diffusion Transformer backbone consistent with prior works [109, 38] and 0.22B new parameters for cross-view attention layers. Training is conducted on 48 NVIDIA A800 GPUs with a batch size of 48. We employ the AdamW optimizer [50] with a learning rate of 1e-4. Further details regarding training objectives, experimental implementations of the occupancy generation module, and the OmniNWM-VLA are provided in the supplementary.
@@ 4.2 Main Results
Table 2: Quantitative evaluation of RGB video generation on the NuScenes validation set. Our proposed method outperforms previous SOTA methods in terms of generation quality.
@@ A.2 Vision-Language Models
Recent years have witnessed growing interest in integrating Vision-Language Models (VLMs) and Large Language Models (LLMs) into autonomous driving systems, capitalizing on their strong reasoning capabilities and world knowledge [2, 104, 46, 59, 15, 85, 66, 107, 33, 108, 97]. These models often exhibit emergent reasoning skills, such as chain-of-thought (CoT) reasoning, which can be harnessed for interpretable decision-making. For example, DriveVLM [66] combines LLM-based trajectory proposals with an end-to-end refinement module for final planning. DriveGPT4 [85] uses LLMs in an iterative QA framework to explain driving behaviors and predict control signals. Doe-1 [107] reformulates driving as a next-token prediction task using a multi-modal generative model, while EMMA [33] encodes non-sensor inputs and outputs as natural language to fully leverage pre-trained LLMs. AutoVLA [111] utilizes raw visual inputs and language instructions for semantic reasoning and trajectory planning. FSDrive [97] introduces a pre-training paradigm for spatio-temporal reasoning, and Impromptu-VLA [11] provides a large-scale annotated dataset and baseline models for vision-language-action tasks in unstructured driving scenarios. In this work, we introduce OmniNWM-VLA, a semantic-geometric reasoning agent built upon Qwen-2.5-VL [2]. It integrates multi-modal inputs—RGB, depth, and semantic maps—via a Mamba-based [20] fusion block, enabling comprehensive scene understanding and trajectory planning.
@@ Appendix D Training Objectives
VLA Planning Objective. The OmniNWM-VLA planner is built upon Qwen-2.5-VL [2] and inherits its training methodology. The model is optimized using standard causal language modeling objective, where the loss is computed as the negative log-likelihood of predicting the next token in the sequence:

### v3

@@ Abstract
Autonomous driving world models are expected to work effectively across three core dimensions: state, action, and reward. Existing models, however, are typically restricted to limited state modalities, short video sequences, imprecise action control, and a lack of reward awareness. In this paper, we introduce OmniNWM, an omniscient panoramic navigation world model that addresses all three dimensions within a unified framework. For state, OmniNWM jointly generates panoramic videos of RGB, semantics, metric depth, and 3D occupancy. A flexible forcing strategy enables high-quality long-horizon auto-regressive generation. For action, we introduce a normalized panoramic Plücker ray-map representation that encodes input trajectories into pixel-level signals, enabling highly precise and generalizable control over panoramic video generation. Regarding reward, we move beyond learning reward functions with external image-based models: instead, we leverage the generated 3D occupancy to directly define rule-based dense rewards for driving compliance and safety. Extensive experiments demonstrate that OmniNWM achieves state-of-the-art performance in video generation, control accuracy, and long-horizon stability, while providing a reliable closed-loop evaluation framework through intrinsic occupancy-grounded rewards.
@@ 1 Introduction
OmniNWM (Ours) R,S,D,O 321 Normalized Plücker Ray-map-based Intrinsic Occupancy-based
@@
  • Leveraging jointly generated 3D occupancy, OmniNWM naturally supports closed-loop navigation and evaluation through intrinsic, occupancy-grounded dense rewards.
@@
Figure 2: Architecture overview of OmniNWM. Our model takes as input a historical trajectory, a reference image, and flexible multi-level noise applied view-wise and frame-wise. It simultaneously forecasts panoramic RGB, semantic, and metric depth videos (Section 3.1.1), along with 3D occupancy for intrinsic dense rewards (Section 3.1.2) and future planning trajectories (Section 3.1.3). Note that input trajectories guide the generation of the current video clip, while future trajectories are forecasted for the next clip, forming a closed-loop system. The input trajectories with camera poses are encoded into normalized panoramic Plücker ray-maps, providing a unified pixel-level description and enabling precise action control (Section 3.2). A flexible forcing strategy injects independent multi-level noise at the view and frame levels to support robust and flexible auto-regressive generation (Section 3.3).
@@ 2 Related Work
World Models for Autonomous Driving. Recent progress in driving world models has advanced visual realism and controllability [100, 18, 116, 42, 16, 44, 79, 8, 98, 90, 19, 99, 66, 109, 58, 34, 36, 49, 64, 45, 76]. Early efforts like DriveDreamer [78] leverage multi-stage pipelines to incorporate traffic rules. MagicDrive [17] introduces cross-view attention for street-view synthesis, while Vista [18] improves generation via latent replacement strategies. UniScene [45] proposes a hierarchical occupancy-centric framework. Other notable works include ReSim [93] for simulation data augmentation, HUGSIM [116] for 3D Gaussian Splatting-based rendering, Drive-WM [79] for multi-view forecasting, and GEM [23] for egocentric prediction. However, existing models typically lack unified multi-modal generation (e.g., semantic map, 3D occupancy)  [78, 17, 18], and integrated reward modeling for closed-loop evaluation [108, 16, 45]. Our work addresses these limitations with long-horizon, multi-modal generation, precise control, and intrinsic rewards.
Camera-controlled Video Generation. Camera control has become a critical technique for guiding generative models in dynamic 3D environments [24, 48, 25, 81, 39, 4, 95, 29]. CameraCtrl [24] introduces a plug-and-play module to parameterize camera trajectories. MotionCtrl [81] decouples camera and object motion for independent control, while CamCo [89] introduces epipolar constraints to improve geometric consistency. VD3D [1] embeds Plücker coordinates into transformer architectures for view synthesis, and ReCamMaster [4] focuses on re-rendering existing videos from novel camera paths. However, these methods focus on short-range, monocular sequences, limiting their applicability to large-scale multi-view environments in autonomous driving [24, 25, 81, 89, 1, 4]. OmniNWM extends camera controllability to panoramic, long-horizon settings through a normalized Plücker ray-map representation.
@@ 3 Methodology
At its core, OmniNWM unifies the state–action–reward triad within a single framework, enabling comprehensive state generation, precise action control, and intrinsic dense rewards. As illustrated in Figure 2, OmniNWM supports multi-modal generation (Section 3.1), producing pixel-aligned panoramic videos (Section 3.1.1), as well as 3D semantic occupancy for intrinsic dense rewards (Section 3.1.2) and future planning trajectories (Section 3.1.3). Section 3.2 introduces the normalized panoramic Plücker ray-map camera control scheme, while Section 3.3 describes the flexible forcing strategy for long-term robust generation.
@@ 3.1 Comprehensive Generation within OmniNWM
The overview of the model architecture is illustrated in Figure 2. We leverage a pretrained 3D VAE [38] to encode input video frames into spatiotemporal latents. A panoramic diffusion transformer (PDiT) with cross-view attention layers [85, 46] is leveraged to denoise the resulting latents. The denoised latents are decoded into pixel-aligned RGB, semantic, and depth outputs, which are subsequently used to produce 3D semantic occupancy and future planning trajectories.
@@ 3.1.1 Pixel-level Aligned Multi-model Generation
where $\otimes$ denotes the outer product. $\texttt{Adap}_{d}$ and $\texttt{Adap}_{s}$ are adaptive aggregation networks for depth and semantic contexts. The resulting volume feature $\widehat{\mathbf{V}}$ is passed to a generation head with up-sampling and a softmax layer to produce the final 3D semantic occupancy. Unlike volumetric-based models like [45, 111] that directly synthesize full occupancy grids at high computational cost, our design computes occupancy from lightweight pixel-aligned features, which scales to long sequences and, crucially, enables intrinsic occupancy-grounded dense driving rewards.
@@ 3.1.3 Trajectory Planning for Closed-loop Navigation
As illustrated in Figure 1(c), OmniNWM functions as an environment simulator to enable closed-loop navigation with planning agents. The planning trajectories generated from the model’s own outputs guide future state generation when encoded into Plücker ray-maps.
@@ 3.2.2 Scale and Pose Invariant Normalization
As shown in Figure 5 (a), we apply scale and pose normalization to each frame ${P}_{j}\in\mathbb{R}^{6\times v\times h\times w}$ of the panoramic ray-map to obtain a unified Plücker representation. Our normalization method decouples camera parameters: a shared intrinsic matrix from a reference camera (defaulting to the front view) ensures scale consistency, while individual extrinsics retain distinct viewpoints. These parameters are transformed into a unified coordinate system. For each pixel $(u,v)$ in the $k^{\text{th}}$ camera, the direction vector is computed using the reference intrinsics ${K}_{0}$ and the camera’s rotation matrix ${R}_{k}$:
@@ 3.3 Long-term Generation with Flexible Forcing
As shown in Figure 6 (b), this formulation enables two distinct auto-regressive inference modes:
(1) Frame-level auto-regression generates a single future frame conditioned on multiple preceding frames:
@@
where $f_{\theta}$ is the diffusion model. $K$ is the input context window. $\mathbf{\tilde{x}}^{(t)}$ is the potentially corrupted latent at time $t$. This mode is particularly suitable for frame-wise trajectory planning simulation (Figure 4).
(2) Clip-level auto-regression generates multiple future frames from fewer inputs:
@@ 4.1 Experimental Setup
We train our model on the NuScenes [6] and NuScenes-Occupancy [77] datasets. Our Panoramic Diffusion Transformer (PDiT) comprises 11.22B parameters, including an 11B Diffusion Transformer backbone consistent with prior works [115, 40] and 0.22B new parameters for cross-view attention layers. Training is conducted on 48 NVIDIA A800 GPUs with a batch size of 48. We employ the AdamW optimizer [55] with a learning rate of 1e-4. Further details regarding training objectives, data processing, experimental implementations of the occupancy generation module, and the OmniNWM-VLA are provided in the supplementary.
@@ 4.2 Main Results
Table 2: Quantitative evaluation of RGB video generation on the NuScenes validation set. Our proposed method outperforms previous SOTA methods without heavy volumetric-based conditions.
@@ A.2 Vision-Language Models
Recent years have witnessed growing interest in integrating Vision-Language Models (VLMs) and Large Language Models (LLMs) into autonomous driving systems, capitalizing on their strong reasoning capabilities and world knowledge [2, 110, 50, 65, 15, 91, 72, 113, 33, 51, 114, 103]. These models often exhibit emergent reasoning skills, such as chain-of-thought (CoT) reasoning, which can be harnessed for interpretable decision-making. For example, DriveVLM [72] combines LLM-based trajectory proposals with an end-to-end refinement module for final planning. SSR [51] addresses the limited spatial understanding of VLMs by converting depth maps into textual rationales and distilling them into lightweight embeddings for plug-and-play spatial reasoning. DriveGPT4 [91] uses LLMs in an iterative QA framework to explain driving behaviors and predict control signals. Doe-1 [113] reformulates driving as a next-token prediction task using a multi-modal generative model, while EMMA [33] encodes non-sensor inputs and outputs as natural language to fully leverage pre-trained LLMs. AutoVLA [118] utilizes raw visual inputs and language instructions for semantic reasoning and trajectory planning. FSDrive [103] introduces a pre-training paradigm for spatio-temporal reasoning, and Impromptu-VLA [11] provides a large-scale annotated dataset and baseline models for vision-language-action tasks in unstructured driving scenarios. In this work, we introduce OmniNWM-VLA, a semantic-geometric reasoning agent built upon Qwen-2.5-VL [2]. It integrates multi-modal inputs—RGB, depth, and semantic maps—via a Mamba-based [20] fusion block, enabling comprehensive scene understanding and trajectory planning.
@@ Appendix D Training Objectives
VLA Planning Objective. The OmniNWM-VLA planner is built upon Qwen-2.5-VL [2] and inherits its training methodology. The model is optimized using the standard causal language modeling objective, where the loss is computed as the negative log-likelihood of predicting the next token in the sequence: